A club with ten members wants to choose a president, vice president, secretary, and tresurer. Six of the members are women, and four are men. What is the probability that all officers will be women?
Please show work?

Answer by jorel1380(3719)

Out of 10 members, 6 are women. The probability that all 4 officers will be women is:
6/10x5/9x4/8x3/7=360/5040=1/14..